# Flow of non-classical correlations in cluster states due to projective measurements

### v1

@@ III.2 Effect of the last measurement
We now focus on the effect of the last measurement on the behavior of non-classical correlations in the qubits on the last run. In our case this measurement is on the $(n-3)^{\rm rd}$ qubit. We keep the values of $\varphi_{i}$, $i=1,\ldots,n-4$ fixed at values that are chosen at random and vary on $\varphi_{n-3}$. Here we compute $D_{n-1\,\bar{n}}$ and $D_{\overline{n-1}\,n}$ as well as the concurrence between the last two qubits for $4$, $6$, $8$ and $10$ qubit ladders. In each case the discord is averaged over all possible measurement outcomes for the preceding measurements. For the first example we consider, the fixed measurement angles chosen for $\varphi_{i}$, $i=1,\ldots,n-4$ (with $\varphi_{n-3}$ varied in the interval 00 to $2\pi$) are given in Table 1:
@@
For the choice of fixed measurement angles in Table 1 we plotted the two discords and concurrence as a function of $\varphi_{n-3}$ in Fig. 8. We see from the plots in Fig. 8 that both $D_{n-1\,\bar{n}}$ and $D_{\overline{n-1}\,n}$ show an increasing trend with increasing number of qubits in the ladder. Furthermore, for the 8 and 10 qubit cases, the last two qubits become entangled as well and the non-classical correlations is progressively dominated by entanglement. The behavior is not strictly monotonic but rather it is so on an average. For instance, $D_{5\bar{6}}$ is zero in the six qubit case for all values of $\varphi_{3}$ even if $D_{3\bar{4}}$ is non-zero for the corresponding 4 qubit case. The concurrence does not show a specific pattern for the two cases (8 and 10 qubits) for which it is non-zero and computations involving more qubits which are prohibitively expensive in terms of compute time will be required to see any trend if it exists.
@@ IV Conclusion
We have explored the non-classical correlations in two-qubit subsystems of ladder-like cluster states when measurements are done on all but the last three qubits. Our results show how projective measurements can generate non-classical correlations in the last two qubits of the ladder-like cluster state. Our study also reveals the flow of non-classical correlations within the cluster state as the measurements are progressively done, throwing light on the dynamics of such correlations during the execution of a quantum algorithm within the MBQC paradigm. We see that with random measurements on $n-3$ qubits in an $n$ qubit ladder, the non-classical correlations between the last two qubits grow stronger on an average when $n$ and the number of rungs of the ladder increases. We also find that with more rungs in the ladder, the last two qubits start to share bipartite entanglement in addition to other more generic non-classical correlations.
For the small clusters with 4 and 6 qubits respectively we were able to understand the behavior of the system under local qubit measurements by computing how the stabilizer generators transform. For both cases, we were able to see that for certain choices of measurement angles, the state of the remaining cluster after the measurements remains a stabiliser state and there will not be any non-classical correlations developed between the two qubits in the last rung. The choices of these measurements are very specific and in almost all cases, the resultant cluster is not a stabilizer state and non-classical correlations are built-up across the last two qubits. This result builds on those in Mahto et al. (2022) wherein a generic connection between multipartite entanglement in structured states like cluster states and sub-system discord was established.

### v2

@@ III.2 Effect of the last measurement
We now focus on the effect of the last measurement on the behaviour of non-classical correlations in the qubits on the last run. In our case this measurement is on the $(n-3)^{\rm rd}$ qubit. We keep the values of $\varphi_{i}$, $i=1,\ldots,n-4$ fixed at values that are chosen at random and vary on $\varphi_{n-3}$. Here we compute $D_{n-1\,\bar{n}}$ and $D_{\overline{n-1}\,n}$ as well as the concurrence between the last two qubits for $4$, $6$, $8$ and $10$ qubit ladders. In each case the discord is averaged over all possible measurement outcomes for the preceding measurements. For the first example we consider, the fixed measurement angles chosen for $\varphi_{i}$, $i=1,\ldots,n-4$ (with $\varphi_{n-3}$ varied in the interval 00 to $2\pi$) are given in Table 1:
@@
For the choice of fixed measurement angles in Table 1 we plotted the two discords and concurrence as a function of $\varphi_{n-3}$ in Fig. 8. We see from the plots in Fig. 8 that both $D_{n-1\,\bar{n}}$ and $D_{\overline{n-1}\,n}$ show an increasing trend with increasing number of qubits in the ladder. Furthermore, for the 8 and 10 qubit cases, the last two qubits become entangled as well and the non-classical correlations is progressively dominated by entanglement. The behaviour is not strictly monotonic but rather it is so on an average. For instance, $D_{5\bar{6}}$ is zero in the six qubit case for all values of $\varphi_{3}$ even if $D_{3\bar{4}}$ is non-zero for the corresponding 4 qubit case. The concurrence does not show a specific pattern for the two cases (8 and 10 qubits) for which it is non-zero and computations involving more qubits which are prohibitively expensive in terms of compute time will be required to see any trend if it exists.
@@
III.3 A single non-Clifford measurement
We look at the dependence of quantum Discord between the qubits on the last rung of the ladder on the location of a specific non-Clifford measurement on one of the remaining $n-3$ qubits. We move this measurements from the first qubit on the first rung of the ladder, step-by-step till the $(n-3)^{\rm rd}$ qubit while performing Clifford measurements on the remaining $n-4$ qubits. Specifically the Clifford measurements done on the $n-4$ qubits are either along the $X$ or $Y$ bases. The results we obtain for ladders with $8$, $10$, $12$ and $14$ qubits respectively are summarized in Fig. 10. in the figure, the Clifford measurements are along the $X$ basis while the non-Clifford measurement on one of the qubits is along $O(\pi/3)$ which corresponds to measurements along $(|0\rangle\pm e^{1\pi/3}|1\rangle)/\sqrt{2}$. We see that the dependence of the quantum discord in the last rung (qubits labelled as $A$ and $B$ and shown in blue in the figure) on the location of the non-Clifford measurement is quite non-intuitive. The locations where the non-Clifford measurement leads to discord between the qubits on the last rung are marked in red in Fig. 10. The values of the discord produced on the last rung are marked near the respective qubits where the non-Clifford measurement is done. The numbers shown are $D^{\rm avg}_{A\bar{B}}$ on the top right, $D^{\rm avg}_{\bar{A}B}$ on the bottom right and the concurrence $C_{AB}$, whenever it is present, on the bottom left of each of the measured qubits. $D^{\rm avg}$ denotes the average value of discord computed over the $2^{n-3}$ possible combinations of “plus” and “minus” measurements on the measured qubits. Non-Clifford measurements on qubits at many locations do not generate any non-classical correlations between $A$ and $B$. These qubits are shown in gray in Fig. 10 and discord values are not listed around the corresponding qubits. It turns out that the values of $D_{A\bar{B}}$ and $D_{\bar{A}B}$ do not depend on whether a “plus” or “minus” projection is done at each site and therefore $D^{\rm avg}_{A\bar{B}}$ and $D^{\rm avg}_{\bar{A}B}$ are just equal to $D_{A\bar{B}}$ and $D_{\bar{A}B}$ for any choice.
  1.01.0\;\;1.01.0  1.01.0\;\;1.01.01.01.01.0\;1.0  0.18870.1887\;\;0.18870.1887  0.18870.1887\;\;0.18870.18870.50.50.5\;0.5  0.18870.1887\;\;0.18870.1887  0.18870.1887\;\;0.18870.18870.50.50.5\;0.5  1.01.0\;\;1.01.0  1.01.0\;\;1.01.01.01.01.0\;1.0  1.01.0\;\;1.01.0  1.01.0\;\;1.01.01.01.01.0\;1.012345CAB 0.16590.1659\;0.16590.1659 0.16590.1659\;0.16590.16591234567CAB 0.16590.1659\;0.16590.1659 0.16590.1659\;0.16590.1659 0.16590.1659\;0.16590.1659 0.16590.1659\;0.16590.1659123456789CAB 0.1650.165\;0.1650.165 0.16590.1659\;0.16590.1659 0.16590.1659\;0.16590.1659 0.16590.1659\;0.16590.16591234567891011CAB
Figure 10: Ladders with different number of rungs are shown along with quantum discord(s) between qubits $A$ and $B$ corresponding to a single non-Clifford measurement $O(\pi/3)$ on one of the $n-3$ other qubits in each ladder. On the remaining $n-4$ qubits, measurements in the $X$ basis is performed. Marked in red are the qubits on which the non-Clifford measurements lead to non-zero discord between $A$ and $B$ (qubits shown in Blue). Non-Clifford measurements on the other qubits shown in gray do not lead to discord between qubits $A$ and $B$. Also shown near the respective qubits are the values of the average discords produced between qubits $A$ and $B$. $D^{\rm avg}_{A\bar{B}}$ is on the top right (blue) and $D^{\rm avg}_{\bar{A}B}$ is on the bottom right (black). The concurrence $C_{AB}$, if present is listed on the bottom left (teal) The average in $D^{\rm avg}$ is over all $2^{n-3}$ combinations of “plus” and “minus” measurements of each of the measured qubits in their respective measurement basis. It turns out that the values of discords are independent of these combinations and they depend only on the location and angle of the single non-Clifford measurement.
We see that there is no discernible pattern for the locations at which non-Clifford measurements lead to discord between the qubits in the last rung. We however notice that the value of $D_{A\bar{B}}$ and $D_{\bar{A}B}$ are independent of the location of the measurement whenever non-zero discord is produced and it depends only on the choice of angle for the non-Clifford measurement. For single non-Clifford measurement corresponding to angles other than $\pi/3$ identical patterns are seen with only the values of $D^{\rm avg}$ changing according to the angle. A similar pattern is seen if the $X$-basis measurements on the remaining $n-4$ qubits are replaced by $Y$-basis measurements. There is no specific pattern available that predicts the location of the non-Clifford measurement that would lead to a finite value of discord. The locations seem quite randomly placed and independent of the number of rungs in the ladder. The ladder with 10 qubits appears to be an exception in that non-Clifford measurements on any one of $n-3$ qubits leads to not only non-zero discord but also non-zero concurrence with the discord and concurrence attaining their maximum possible values in several cases. This again serves to highlight the non intuitive manner in which quantum information distributes itself across the qubits in the ladder and how it is transformed under measurement on all but the last three qubits.
@@ IV Conclusion
We have explored the non-classical correlations in two-qubit subsystems of ladder-like cluster states when measurements are done on all but the last three qubits. Our results show how projective measurements can generate non-classical correlations in the last two qubits of the ladder-like cluster state. Our study also reveals the flow of non-classical correlations within the cluster state as the measurements are progressively done, throwing light on the dynamics of such correlations during the execution of a quantum algorithm within the MBQC paradigm. We see that with random measurements on $n-3$ qubits in an $n$ qubit ladder, the non-classical correlations between the last two qubits grow stronger on an average when $n$ and the number of rungs of the ladder increases. We also find that with more rungs in the ladder, the last two qubits start to share bipartite entanglement in addition to other more generic non-classical correlations. With only a single non-Clifford measurement, we find that the flow of correlations through the ladder is quite non-intuitive, highlighting the different and unexpected ways in which delocalized quantum information in the ladder can be modulated and moved around through measurements on various qubits on the rungs of the ladder.
For the small clusters with 4 and 6 qubits respectively we were able to understand the behavior of the system under local qubit measurements by computing how the stabilizer generators transform. For both cases, we were able to see that for certain choices of measurement angles, the state of the remaining cluster after the measurements remains a stabiliser state and there will not be any non-classical correlations developed between the two qubits in the last rung. The choices of these measurements are very specific and in almost all cases, the resultant cluster is not a stabilizer state and non-classical correlations are built-up across the last two qubits. This result builds on those in [25] wherein a generic connection between multipartite entanglement in structured states like cluster states and subsystem discord was established.